OpenStudy (anonymous):

What is the solution to this system of equations? X=10 Y=2x-6 A.no solutions B.(10,14) C.(20,10) D.(2,6)

OpenStudy (isaidavila):

It's 10 and 14 just replace the value of x un y an that's all

OpenStudy (richgall____):

@isaidavila i understand the 10 part but where do you get 14 from ?

OpenStudy (richgall____):

nvm i see it

OpenStudy (richgall____):

y = 2x - 6 y = (2 x 10) - 6 y = 20 - 6 20 - 6 = 14

OpenStudy (isaidavila):

Yes replace x in y